# Interact2Vec – An efficient neural network-based model for simultaneously learning users and items embeddings in recommender systems

### v1

@@ Abstract
[…]
@@
Copyright ©  [2025] Elsevier. This is a post-peer-review version of an article to be published in Applied Soft Computing. This manuscript is made available under the Elsevier user license.
Accepted to be published in: Applied Soft Computing, 2025.
@@
Pedro R. Pires, and Tiago A. Almeida. (2025). “Interact2Vec – An efficient neural network-based model for simultaneously learning users and items embeddings in recommender systems,” in Applied Soft Computing.
@@
     @article{Pires2025,    author    = {Pedro. R. Pires and Tiago A. Almeida},    title     = {Interact2Vec -- An efficient neural network-based                   model for simultaneously learning users and items                   embeddings in recommender systems},    journal   = {Applied Soft Computing},    publisher = {Elsevier B.V.},    year      = {2025},  }
@@ 4.1 Datasets and preprocessing
Anime [^1] Anime Recommendations dataset. Available at: https://www.kaggle.com/datasets/CooperUnion/anime-recommendations-database. Accessed on June 27, 2025. 37,128 10,697 1,476,495 99.63% E+I
BestBuy [^2] Data Mining Hackathon on BIG DATA (7GB). Available at: www.kaggle.com/c/acm-sf-chapter-hackathon-big. Accessed on June 27, 2025. 1,268,702 69,858 1,862,782 99.99% I
BookCrossing [^3] GroupLens - Datasets. Available at: www.grouplens.org/datasets/. Accessed on June 27, 2025. 59,517 246,724 716,109 99.99% E+I
CiaoDVD [^4] CiaoDVD dataset and Filmtrust. Available at: www.github.com/caserec/Datasets-for-Recommender-Systems. Accessed on June 27, 2025. 17,615 16,121 72,345 99.97% E
Delicious [^3] GroupLens - Datasets. Available at: www.grouplens.org/datasets/. Accessed on June 27, 2025. 1,867 69,223 104,799 99.92% I
Filmtrust [^4] CiaoDVD dataset and Filmtrust. Available at: www.github.com/caserec/Datasets-for-Recommender-Systems. Accessed on June 27, 2025. 1,508 2,071 35,494 98.86% E
Last.FM [^3] GroupLens - Datasets. Available at: www.grouplens.org/datasets/. Accessed on June 27, 2025. 1,892 17,632 92,834 99.72% I
MovieLens [^3] GroupLens - Datasets. Available at: www.grouplens.org/datasets/. Accessed on June 27, 2025. 162,541 59,047 25,000,095 99.74% E
NetflixPrize [^5] Netflix Prize data. Available at: www.kaggle.com/netflix-inc/netflix-prize-data. Accessed on June 27, 2025. 480,189 17,770 100,480,507 98.82% E
RetailRocket [^6] Retailrocket recommender system dataset. Available at www.kaggle.com/retailrocket/ecommerce-dataset. Accessed on June 27, 2025. 11,719 12,025 21,270 99.98% I
@@ 4.2 Experimental protocol
All algorithms were implemented in Python3. For KNN we used the implementation of turicreate library. Both ALS and BPR were implemented using implicit library, with latent factors $f$ ranging in $\{50,100,200\}$, regularization factor $\lambda$ in $\{10^{-6},10^{-4},10^{-2}\}$ and learning rate $\alpha$ in $\{0.0025,0.025,0.25\}$. IT2V and US2V were implemented using gensim library, with $C$ ranging in $\{50,100,150\}$, $\gamma$ in $\{-1.0,-0.5,0.5,1.0\}$, and $\rho$ in $\{10^{-5},10^{-4},10^{-3}\}$, while the remaining hyperparameters followed insights proposed in [41], with fixed optimal values for the parameters shown in Table 3. Lastly, RecVAE (RVAE) was tested with its original implementation [^2] RecVAE source code. Available at: https://github.com/ilya-shenbin/RecVAE. Accessed on June 27, 2025., with $\gamma$ ranging in $\{0.0035,0.005,0.01\}$ and the remaining hyperparameters fixed to the values recommended by its authors [45].

### v2

@@ Abstract
Over the past decade, recommender systems have experienced a surge in popularity. Despite notable progress, they grapple with challenging issues, such as high data dimensionality and sparseness. Representing users and items as low-dimensional embeddings learned via neural networks has become a leading solution. However, while recent studies show promising results, many approaches rely on complex architectures or require content data, which may not always be available. This paper presents Interact2Vec, a novel neural network-based model that simultaneously learns distributed embeddings for users and items while demanding only implicit feedback. The model employs state-of-the-art strategies that natural language processing models commonly use to optimize the training phase and enhance the final embeddings. Two types of experiments were conducted regarding the extrinsic and intrinsic quality of the model. In the former, we benchmarked the recommendations generated by Interact2Vec’s embeddings in a top-$N$ ranking problem, comparing them with six other recommender algorithms. The model achieved the second or third-best results in 30% of the datasets, being competitive with other recommenders, and has proven to be very efficient with an average training time reduction of 274% compared to other embedding-based models. Later, we analyzed the intrinsic quality of the embeddings through similarity tables. Our findings suggest that Interact2Vec can achieve promising results, especially on the extrinsic task, and is an excellent embedding-generator model for scenarios of scarce computing resources, enabling the learning of item and user embeddings simultaneously and efficiently.
@@
Copyright ©  [2025] Elsevier. This is a post-peer-review version of an article published in Applied Soft Computing. This manuscript is made available under the Elsevier user license.
Published in: Applied Soft Computing, 2025.
@@
Pedro R. Pires, Tiago A. Almeida, Interact2Vec – An efficient neural network-based model for simultaneously learning users and items embeddings in recommender systems, Applied Soft Computing, 181 (2025) 113408:1–113408:17, http://dx.doi.org/ 10.1016/j.asoc.2025.113408.
@@
     @article{PIRES2025113408,      title    = {Interact2Vec -- An efficient neural network-based                  model for simultaneously learning users and items                  embeddings in recommender systems},      journal  = {Applied Soft Computing},      volume   = {181},      pages    = {113408:1--113408:17},      year     = {2025},      issn     = {1568-4946},      doi      = {https://doi.org/10.1016/j.asoc.2025.113408},      url      = {https://www.sciencedirect.com/science/article/pii/                  S1568494625007197},      author   = {Pedro R. Pires and Tiago A. Almeida},      keywords = {Recommender systems, Collaborative filtering,                  Distributed vector representation, Embeddings},  }
@@ 4.1 Datasets and preprocessing
Anime [^1] Anime Recommendations dataset. Available at: https://www.kaggle.com/datasets/CooperUnion/anime-recommendations-database. Accessed on July 2, 2025. 37,128 10,697 1,476,495 99.63% E+I
BestBuy [^2] Data Mining Hackathon on BIG DATA (7GB). Available at: www.kaggle.com/c/acm-sf-chapter-hackathon-big. Accessed on July 2, 2025. 1,268,702 69,858 1,862,782 99.99% I
BookCrossing [^3] GroupLens - Datasets. Available at: www.grouplens.org/datasets/. Accessed on July 2, 2025. 59,517 246,724 716,109 99.99% E+I
CiaoDVD [^4] CiaoDVD dataset and Filmtrust. Available at: www.github.com/caserec/Datasets-for-Recommender-Systems. Accessed on July 2, 2025. 17,615 16,121 72,345 99.97% E
Delicious [^3] GroupLens - Datasets. Available at: www.grouplens.org/datasets/. Accessed on July 2, 2025. 1,867 69,223 104,799 99.92% I
Filmtrust [^4] CiaoDVD dataset and Filmtrust. Available at: www.github.com/caserec/Datasets-for-Recommender-Systems. Accessed on July 2, 2025. 1,508 2,071 35,494 98.86% E
Last.FM [^3] GroupLens - Datasets. Available at: www.grouplens.org/datasets/. Accessed on July 2, 2025. 1,892 17,632 92,834 99.72% I
MovieLens [^3] GroupLens - Datasets. Available at: www.grouplens.org/datasets/. Accessed on July 2, 2025. 162,541 59,047 25,000,095 99.74% E
NetflixPrize [^5] Netflix Prize data. Available at: www.kaggle.com/netflix-inc/netflix-prize-data. Accessed on July 2, 2025. 480,189 17,770 100,480,507 98.82% E
RetailRocket [^6] Retailrocket recommender system dataset. Available at www.kaggle.com/retailrocket/ecommerce-dataset. Accessed on July 2, 2025. 11,719 12,025 21,270 99.98% I
@@ 4.2 Experimental protocol
All algorithms were implemented in Python3. For KNN we used the implementation of turicreate library. Both ALS and BPR were implemented using implicit library, with latent factors $f$ ranging in $\{50,100,200\}$, regularization factor $\lambda$ in $\{10^{-6},10^{-4},10^{-2}\}$ and learning rate $\alpha$ in $\{0.0025,0.025,0.25\}$. IT2V and US2V were implemented using gensim library, with $C$ ranging in $\{50,100,150\}$, $\gamma$ in $\{-1.0,-0.5,0.5,1.0\}$, and $\rho$ in $\{10^{-5},10^{-4},10^{-3}\}$, while the remaining hyperparameters followed insights proposed in [41], with fixed optimal values for the parameters shown in Table 3. Lastly, RecVAE (RVAE) was tested with its original implementation [^2] RecVAE source code. Available at: https://github.com/ilya-shenbin/RecVAE. Accessed on July 2, 2025., with $\gamma$ ranging in $\{0.0035,0.005,0.01\}$ and the remaining hyperparameters fixed to the values recommended by its authors [45].